OpenStudy (anonymous):

use Gauss's approach to find the following sums (do not use formulas). a. 1+2+3+4+...+98 b. 1+3+5+7+...+997 a. the sum of the sequence is?

OpenStudy (anonymous):

after solving for a it will want u to solve for b i dont understand it at all though.

OpenStudy (anonymous):

isnt this where you just reverse the order and add them together?\[Sum=1+2+\cdots +97+98\]\[Sum=98+97+\cdots+2+1\]adding them gives: \[2\cdot Sum=(1+98)+(2+97)+\cdots + (97+2)+(98+1)\]\[\iff 2\cdot Sum = \underbrace{99+99+\cdots +99+99}_{\frac{98}{2} \mbox \times}\] \[\iff 2\cdot Sum = 99\cdot 98\iff Sum = \frac {99\cdot 98}{2}\]

OpenStudy (anonymous):

oops, that should be "98 times" under the bracket.

OpenStudy (anonymous):

so is this easy but they make it look complicateted?

OpenStudy (anonymous):

i guess o.O its easy for some but not easy for all.

OpenStudy (anonymous):

how would i solve for the second one?